A hydrocarbon gas has a density of 1.18 g/L at 20.°C and 1.00 atm. An analysis gives 85.7% C and 14.3% H. What is the molecular formula?
Answer #1

density (d) = 1,18 g/L Temperature (T) = 20°C = 20 + 273k = 293k pressure (P)= I atm. So from ideal gas d = pm m (molecular mMolecular formula = (empirical formalan = (CH2), санч